# A Super Version of a theorem of Fricke-Klein

### v1

@@ Abstract.
In this paper, we study the character variety of the supergroup $\operatorname{OSp}(1|2)$ in the specific case of the free group on two letters as a first step towards an algebraic treatment for the character variety of supergroups. One can interpret this as an analogue of a result of Fricke and Klein, who studied this in the case of $\operatorname{SL}(2,\mathbb{C})$. We give a count of independent generators of the ring of invariants.
@@ 1. Introduction
A result by Fricke and Klein [8] and apparently Vogt [25] states, that the $SL(2,\mathbb{C})$-invariant functions on $SL(2,\mathbb{C})\times SL(2,\mathbb{C})$ are given by the ring
@@
In the context of geometric invariant theory, we can interpret this result as
$\operatorname{Spec}(SL_{2})\times\operatorname{Spec}(SL_{2})$$\operatorname{Spec}(\mathbb{C}[\operatorname{tr}(A),\operatorname{tr}(B),\operatorname{tr}(AB)])\cong\mathbb{C}^{3}$$\scriptstyle{\pi}$
i.e. a quotient variety. Similarly, another reinterpretation of this result , which is even more geometric, is that it is equivalent to the fact that the character variety of the free group on two letters and $SL(2,\mathbb{C})$ is isomorphic to $\mathbb{C}^{3}$, i.e.
@@
In this paper we are aiming to generalize the third interpretation of this result, to the super world. The motivation of this result comes from the fact, that $\operatorname{OSp}(1|2)$ is the super generalization of $\operatorname{SL}(2,\mathbb{C})$ and takes a similar place in the theory of super Riemann surfaces, as $\operatorname{SL}(2,\mathbb{C})$ does in the classical theory. On the real analytic side of this story, we get something called super Teichmüller space, where each connected component of the space is homeomorphic to $\mathbb{R}^{6g-6|4g-4}/\mathbb{Z}_{2}$ [2] [19], where $\mathbb{Z}_{2}$ only acts on the odd coordinates.
By ”explicit” computation, we can show that there is a deviation from the classical case and we do not get a variety of the expected dimension $3|2$. A problem of supergeometry is that most of the time the super version of the proof works analogously without adjusting anything. We want to close the gap of the unknown by writing down the super versions of the proof, even if nothing changes, so the reader does not have to mentally go through the classical construction and reason, that it works accordingly in the super case.
Organization of the paper. In section 2 we will give a short exposition on the result of Fricke and Klein. In section 3 introduce basic notions of supergeometry, algebraic supergroups and invariant theory to understand our main result. In section 4 we introduce the character variety, which is the geometric interpretation of our result, where we can also easily see that this definition adapts to super. In section 5 we prove our main result
@@
The $\operatorname{OSp}(V)$-invariant polynomial functions (by conjugation) on $\operatorname{OSp}(V)\times\operatorname{OSp}(V)$ are generated by 7 traces.
Acknowledgments. I would like to thank my advisor Paul Norbury for his constant guidance and confidence in my work. I would also like to thank Miles Koumouris, Gustav Berth and Lior Yanovski for useful discussion.
@@ 2. Fricke-Klein
In this section, we give an elementary proof of the result of Fricke and Klein and apparently Vogt. In more recent times, Goldman reproved this theorem in [10], where he used the Cayley-Hamilton theorem as in an essential way in the proof. In this version, we forego the usage of the Cayley-Hamilton theorem and instead make use of a simultaneous conjugation of a pair of matrices into upper and lower triangular forms. Before we start proving the result of Fricke and Klein, let us first state the aforementioned triangulation lemma.
@@
(3) $gA_{0}g^{-1}=\begin{pmatrix}\mu&0\\ 1&\mu^{-1}\end{pmatrix}\quad,\quad B_{0}=\begin{pmatrix}\lambda&\kappa\\ 0&\lambda^{-1}\end{pmatrix}$
@@
This splits into two cases. $A$ diagonalizable or not. If $A$ is diagonalizable, then it must mean that for its eigenvalues $\mu,\mu^{-1}$ that $\mu\neq\mu^{-1}$. Consider
@@
Let $A$ not diagonalizable, then it must admit Jordan normal form
@@
$A\in\operatorname{SL}(2,\mathbb{C})$, therefore $\mu=1$, which is a case we can treat by hands. Consider
$\begin{pmatrix}x&y\\ 0&z\end{pmatrix}\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\begin{pmatrix}z&-y\\ 0&x\end{pmatrix}=\begin{pmatrix}1&x^{2}\\ 0&1\end{pmatrix}$
@@
Also notice, that
@@
First, we set $x=i$, then we choose a solution for the equation y(xa+yc)+x(xb+yd)-y(xa+yc)+x(xb+yd), which we can do over $\mathbb{C}$. Act by rotation by 90 degree matrix. and we get the pair
$A=\begin{pmatrix}1&0\\ 1&1\end{pmatrix}\quad,\quad B=\begin{pmatrix}z(xa+yc)&z^{2}c\\ 0&-yzc+xzd\end{pmatrix}$,
which is in the desired form. So, indeed we can always parametrize into the simultaneous triangular form. ∎
@@
where $X=\operatorname{tr}(A)$, $Y=\operatorname{tr}(B)$ and $Z=\operatorname{tr}(AB)$. First of all note, that if $f\in\mathbb{C}[H]$, then $f$ is a polynomial function in the entries of the two matrices $(\xi,\eta)\in SL(2,\mathbb{C})\times SL(2,\mathbb{C})$. We will denote it $f(A,B)$. Furthermore, by the triangulation lemma we can bring them into the form
@@
So it turns out, that the statement is even stronger, we have
@@
which is surjective as argued earlier. Now to prove it is injective, we again make use of our identifications
@@
$\sum_{ijk}\alpha_{ijk}\sum_{l}^{i}\sum_{m}^{j}\sum_{k_{1}+k_{2}=k_{3}=n}^{k}\binom{i}{l}\binom{j}{m}\binom{k}{k_{1},k_{2},k_{3}}\lambda^{i-2l+k_{1}-k_{2}}\mu^{j-2m+k_{1}-k_{2}}\kappa^{k_{3}}$
On the other hand we have the $G$-invariant polynomial
@@
We can clearly see, every $\beta_{rst}$ is just a sum of the $\alpha_{ijk}$ By our condition we know all coefficients $\alpha_{ijk}$ must vanish, since there is no relation between $X,Y,Z$. Therefore $\beta_{rst}=0$. Thus $g=0$ and injectivity follows. In total we now have
$\mathbb{C}[G\times G]^{G}=\mathbb{C}[H]\cong\mathbb{C}[X,Y,Z]$
@@
$\operatorname{Hom}(F_{2},SL(2,\mathbb{C}))//SL(2,\mathbb{C})\cong\mathbb{C}^{3}$
@@ 3.1. Supergeometry
Here we want to introduce basic notions of supergeometry following [11] and [3]. Generally, we will simply endow everything with a $\mathbb{Z}_{2}$-grading.
@@
Let $V=V_{0}\oplus V_{1}$ be a $\mathbb{Z}_{2}$-graded vector space, then it is a super vector space
@@
A superscheme $X=(|X|,\mathcal{O}_{X})$ is a locally ringed space, where $|X|$ is a topological space and $\mathcal{O}_{X}$ is a sheaf of $\mathbb{Z}_{2}$-graded commutative rings.
@@
We can interpret the odd coordinates geometrically as fat points. To elaborate a little bit on that:
Odd coordinates vanish if you square them, so if we think of a scheme over $\mathbb{C}$ of odd dimension $0|1$, we can take a look at the coordinate ring $\mathbb{C}[\theta]$ which is roughly the same as $\mathbb{C}[\epsilon]/\epsilon^{2}$, where $\theta$ is an odd coordinate, while $\epsilon$ is an even one.
By this analogy, we can think of odd coordinates geometrically as smearing out the point or a first order taylor expansion around the origin.
Example 3.4.
$\mathbb{C}^{n|m}$.
Example 3.5.
$\mathbb{P}^{n|m}$. Here we want to construct the super version of projective space. As underlying topological space we use $\mathbb{P}^{n}$, which is a collection of $U_{i}=\{[z_{0}:...:z_{n}]|z_{i}\neq 0\}$. These subsets are isomorphic to $\mathbb{C}^{n+1}$, because $z_{i}$ is always non zero, so we can divide by $z_{i}$ and get $U_{i}=\{[\frac{z_{0}}{z_{i}}:,...,1,...,\frac{z_{n}}{z_{i}}]\}$ . So if we take the corresponding locally free sheaf of rank m, then the associated supermanifold $(U_{i},\mathcal{O}_{U_{i}})$ is going to be $\mathbb{C}^{n|m}$. We can also define an isomorphism on the set $U_{i}\cap U_{j}$. We can choose n even generating sections of $U_{i}$, called $u_{i,a}$, and m odd generating sections called $\xi_{i,b}$. So we write:
(5) $\phi_{ij}(u_{i,0},...,\hat{u}_{i,i},...,u_{i,n},\xi_{i,1},...,\xi_{i,m})\mapsto(\frac{u_{i,0}}{u_{i,j}},..,\frac{1}{u_{i}},...,\frac{u_{i,n}}{u_{i,j}},\frac{\xi_{i,1}}{u_{i,j}},...,\frac{\xi_{i,m}}{u_{i,j}})$
We can also do this the other way around and can see, that the transition function on the intersection is $\frac{u_{i,i}}{u_{i,j}}$. One can also easily see, that these $\phi_{ij}$ satisfy the cocycle condition. So we can glue all these sections together and define a sheaf $\mathcal{O}_{\mathbb{P}^{n|m}}$. Together with the underlying topological space we have defined the super projective space $(\mathbb{P}^{n},\mathcal{O}_{\mathbb{P}^{n|m}})$.
@@
The reduced space of a superscheme $X$ is defined as
(6) $X\coloneqq(|X|,\mathcal{O}_{X}/J_{X})=(|X|,\mathcal{O}_{red})$
where $J_{X}$ is the sheaf of ideals generated by the nilpotent sections of $\mathcal{O}_{X}$.
It is important to note, that $\mathcal{O}_{X}/J\coloneqq a(U\mapsto\mathcal{O}_{X}(U)/J(U))$ is the sheafification. However, we still get an isomorphism on stalks between the presheaf and its sheafification and hence we can use results known from commutative algebra with impunity. This notion is useful, because when we try to generalize our concepts to try to describe more abstract physics, we want to ensure that it does reduce to our known description of phenomena up to that point. However, there is generally an obstruction to being able to do this nicely. Ideally, we want a global section to exist, which is the projection onto the non-super space. So, we will introduce the notion of projected and split superscheme.
Definition 3.7.
A superscheme X is called projected if the structural sequence
@@
As one might guess, this implies the existence of non-projected superscheme, but contrary to what our intuition might lead us to conclude a lot of supermanifolds are non-projected in even quite simple examples.
@@
In the category of supermanifolds, it turns out that every real supermanifold is split, due to Batchelors theorem (in a non-natural way). So if we want to look for non-projected objects, we must look at complex supermanifolds.
@@
Let X be a superscheme of odd dimension m, then there exists a $J_{X}$ adic filtration of $\mathcal{O}_{X}$
@@
(8) $Gr(\mathcal{O}_{X})\coloneqq J_{X}^{0}/J_{X}^{1}\oplus J_{X}^{1}/J_{X}^{2}\oplus...\oplus J_{X}^{m-1}/J_{X}^{m}\oplus J^{m}_{X}$
The split superscheme associated to X is defined as
(9) $Gr(X)\coloneqq(|X|,Gr(\mathcal{O}_{X})).$
@@
A superscheme $X$ is called split, if there exists an isomorphism $X\cong Gr(X)$.
Lemma 3.11.
If a supermanifold X of odd dimension $m$ is split, it is projected.
@@
The terminology in supergeometry is not yet set in stone. Some authors refer to this property as split and some authors call this property being a graded supermanifold. We remind the reader to always carefully check, what is meant by the term graded or split when encountering supergeometry.
Remark 3.13.
A more categorical way and maybe more uniform way to view supergeometry in the context of ”prefix” algebraic geometry. Let $\operatorname{Ab}_{\mathbb{Z}_{2}}$ be the category of $\mathbb{Z}_{2}$-graded abelian groups, which is symmetric monoidal, then a superring $A$ is an commutative algebra/commutative monoid object inside of $\operatorname{Ab}_{\mathbb{Z}_{2}}$, which is usually denoted $\operatorname{CAlg}(\operatorname{Ab}_{\mathbb{Z}_{2}})\cong\operatorname{SRings}$.
3.2. Supergroups
Here we will introduce algebraic supergroups following [16] [15], the analogue of algebraic groups in supergeometry. Algebraic supergroups naturally appear, when dealing with supersymmetry related models in physics, however a more mathematics oriented answer as in how they appear is that representations of supergroups naturally appear in Delignes theorem on tensor categories [6]. One heuristic way of why they appear is that the category of vector spaces is not algebraically closed so to say, there is a natural $\mathbb{Z}_{2}$-action on the category of vector spaces and taking its closure with respect to that will result in the category of super vector spaces. So when you are trying to turn a tensor category into a category of representations you will see that if you restrict to ordinary vector spaces that you simply do not have enough points in usual setting one would move from $\mathbb{R}$ to $\mathbb{C}$, while in this case on moves from $\mathrm{Vect}_{k}$ to $\mathrm{sVect}_{k}$ for $k$ algebraically closed and of characteristic 0.
In standard literature schemes are described as locally ringed spaces equipped with the appropriate sheaf, that extracts the algebraic structure. This formalism allows us to easily enter the realm of supergeometry, by replacing our sheaves with $\mathbb{Z}_{2}$-graded sheafs. Similarly, algebraic groups can be studied as a locally ringed space. In the affine case, to make the group structure compatible/easily translatable through our functor Spec, we should describe it as a set of arrows that we can invert, when applying Spec. This procedure results in the following

### v2

@@ Abstract.
In this paper, we study the character variety of the supergroup $\operatorname{OSp}(1|2)$ in the specific case of the free group on two letters as a first step towards an algebraic treatment for the character variety of supergroups. One can interpret this as an analogue of a result of Fricke and Klein, who studied this in the case of $\operatorname{SL}(2,\mathbb{C})$. We give a count of independent generators of the ring of invariants and discuss a few geometric consequences at the end.
@@ 1. Introduction
A result by Fricke and Klein [10] states, that the $SL(2,\mathbb{C})$-invariant functions on $SL(2,\mathbb{C})\times SL(2,\mathbb{C})$ are given by the ring
@@
In the language of geometric invariant theory, we can identify the categorical quotient of the product variety $\mathrm{SL}_{2}(\mathbb{C})\times\mathrm{SL}_{2}(\mathbb{C})$ by conjugation as affine space, i.e.
$\mathrm{SL}_{2}(\mathbb{C})\times\mathrm{SL}_{2}(\mathbb{C})$$\operatorname{Spec}(\mathbb{C}[\operatorname{tr}(A),\operatorname{tr}(B),\operatorname{tr}(AB)])\cong\mathbb{C}^{3}.$$\scriptstyle{\pi}$
Another geometric reinterpreation of this result is that the character variety of the free group on two letters and $SL(2,\mathbb{C})$ is isomorphic to $\mathbb{C}^{3}$, i.e.
@@
In this paper we establish a variant of the theorem of Fricke and Klein for a supergeometric extension $\mathrm{SL}_{2}(\mathbb{C})$, specifically the complex supergroup $\mathrm{OSp}(1|2)$. The motivation of this result comes from the fact, that $\operatorname{OSp}(1|2)$ is the super generalization of $\operatorname{SL}(2,\mathbb{C})$ and takes a similar place in the theory of super Riemann surfaces, as $\operatorname{SL}(2,\mathbb{C})$ does in the classical theory. On the real analytic side of this story, the character variety is related to the super Teichmüller space, where each connected component of the space is homeomorphic to $\mathbb{R}^{6g-6|4g-4}/\mathbb{Z}_{2}$ [3] [25], where $\mathbb{Z}_{2}$ only acts on the odd coordinates.
We will show that the character variety for the free group on two letters $F_{2}$ and the supergroup $\mathrm{OSp}(1|2)$ fails to be a superscheme. Many results in supergeometry can be obtained by the same proof as in the classical setting and the literature frequently leaves the explicit verification to the reader. In this work we opt for a self-contained exposition for the necessary results adapted to the super setting. Organization of the paper. In section 2 we will give a short exposition on the result of Fricke and Klein. In section 3, we start by introducing basic notions of supergeometry, algebraic supergroups and invariant theory to understand our main result. In section 4 we prove our main result. In section 5 we introduce the character variety and the character stack, which is the geometric interpretation of our result, and discuss the geometric implication of the main result.
@@
The $\operatorname{OSp}(V)$-invariant polynomial functions (by conjugation) on $\operatorname{OSp}(V)\times\operatorname{OSp}(V)$ are generated by 7 independent traces.
Acknowledgments. I would like to thank my advisor Paul Norbury for his constant guidance and confidence in my work. I would also like to thank Gustav Berth, Miles Koumouris and Lior Yanovski for useful discussion.
@@ 2. Fricke-Klein
In this section, we give an elementary proof of the result of Fricke and Klein [10] and separately Vogt [33]. In more recent times, Goldman reproved this theorem in [12], where he used the Cayley-Hamilton theorem in an essential way in the proof. In this version, we forego the usage of the Cayley-Hamilton theorem and instead make use of a simultaneous conjugation of a pair of matrices into upper and lower triangular forms. Before we start proving the result of Fricke and Klein, let us first state the aforementioned triangulation lemma.
@@
(3) $A_{0}\coloneqq gAg^{-1}=\begin{pmatrix}\mu&0\\ 1&\mu^{-1}\end{pmatrix}\quad,\quad B_{0}\coloneqq gBg^{-1}=\begin{pmatrix}\lambda&\kappa\\ 0&\lambda^{-1}\end{pmatrix}$
@@
This splits into two cases. $A$ diagonalizable or not. If $A$ is diagonalizable, we assume that $A=\mathrm{diag(\mu,\mu^{-1})}$, such that $\mu\neq\mu^{-1}$. Consider
@@
If $A$ not diagonalizable, then it must admit Jordan normal form
@@
As $A\in\operatorname{SL}(2,\mathbb{C})$, we have $\mu=\pm 1$, which is a case we can treat by hand. Consider
$\begin{pmatrix}x&y\\ 0&x^{-1}\end{pmatrix}\begin{pmatrix}\pm 1&1\\ 0&\pm 1\end{pmatrix}\begin{pmatrix}x^{-1}&-y\\ 0&x\end{pmatrix}=\begin{pmatrix}\pm 1&x^{2}\\ 0&\pm 1\end{pmatrix}$
@@
Also note that the reflection matrix acts on upper/lower triangular matrices as
@@
and
$\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\begin{pmatrix}a&0\\ c&d\end{pmatrix}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}=\begin{pmatrix}a&-c\\ 0&d\end{pmatrix}$
First, we set $x=i$, then we choose a solution for the equation y(xa+yc)+x(xb+yd)-y(xa+yc)+x(xb+yd), which we can do over $\mathbb{C}$. We multiply by rotation matrix, which corresponds to rotating by $\frac{\pi}{2}$. Thus, we obtain the pair
$A=\begin{pmatrix}1&0\\ 1&1\end{pmatrix}\quad,\quad B=\begin{pmatrix}z(xa+yc)&-z^{2}c\\ 0&-yzc+xzd\end{pmatrix}$,
which is in the desired form. ∎
@@
where $X=\operatorname{tr}(A)$, $Y=\operatorname{tr}(B)$ and $Z=\operatorname{tr}(AB)$. First of all note that if $f\in\mathbb{C}[H]$, then $f$ is a polynomial function in the entries of the two matrices $(\xi,\eta)\in SL(2,\mathbb{C})\times SL(2,\mathbb{C})$. We will denote it $f(A,B)$. Furthermore, by the triangulation lemma we can bring them into the form
@@
We observe that the statement is even stronger, we have
@@
which is surjective as argued earlier. We now prove it is injective. We again make use of our identifications
@@
$\sum_{ijk}\alpha_{ijk}\sum_{l}^{i}\sum_{m}^{j}\sum_{k_{1}+k_{2}+k_{3}=n}^{k}\binom{i}{l}\binom{j}{m}\binom{k}{k_{1},k_{2},k_{3}}\lambda^{i-2l+k_{1}-k_{2}}\mu^{j-2m+k_{1}-k_{2}}\kappa^{k_{3}}$
On the other hand, we have the $G$-invariant polynomial
@@
We observe that every $\beta_{rst}$ is a sum of the $\alpha_{ijk}$ By our condition, we know that all coefficients $\alpha_{ijk}$ must vanish, since there is no relation between $X,Y,Z$. Therefore $\beta_{rst}=0$. Thus $g=0$ and injectivity follows. In total we now have
$\mathbb{C}[G\times G]^{G}=\mathbb{C}[H]\cong\mathbb{C}[X,Y,Z],$
@@
$\operatorname{Hom}(F_{2},SL(2,\mathbb{C}))//SL(2,\mathbb{C})\cong\mathbb{C}^{3}.$
@@ 3.1. Supergeometry
Here we want to introduce basic notions of supergeometry. In particular, we will provide a categorical viewpoint, alongside its classical definitions, thereby highlighting the similarities between commutative algebras and supercommutative algebra. This should aid the reader to understand which phenomena will directly carry over and in which situations some more care ought to be given. A possible way to view the category of commutative rings is that it is the category of commutative monoid objects with respect to the tensor product in the symmetric monoidal category $\mathrm{Ab}$. To do supercommutative algebra and eventually supergeometry, the only shift that will occur from the categorical perspective is that we replace the category $\mathrm{Ab}$ with $\mathrm{Ab}_{\mathbb{Z}_{2}}$. In practice, all objects will be endowed with a $\mathbb{Z}_{2}$-grading and a Koszul sign rule. Unless otherwise note, the definitions in this section follow [4].
@@
Let $\mathrm{Ab}_{\mathbb{Z}_{2}}$ be the category of abelian groups with a $\mathbb{Z}_{2}$-grading, which we equip with the symmetric monoidal structure given by the braiding isomorphism
(5) $\begin{split}M\otimes N&\longrightarrow N\otimes M\\ m\otimes n&\longmapsto(-1)^{|m||n|}n\otimes m,\end{split}$
where $|m|$ denotes the degree of the element.
Remark 3.2.
We will be using both $\mathrm{deg}(-)$ and $|\cdot|$ to denote the degree map.
Definition 3.3 ([35]).
A superring $A$ is a $\mathbb{Z}_{2}$-graded ring $A=A_{0}\oplus A_{1}$ such that the product map $A\times A\to A$ satisfies $A_{i}A_{j}\subset A_{i+j}$. A morphism of superrings is a $\mathbb{Z}_{2}$-grading preserving morphism of rings. The elements of $A_{0}$ are called even, the elements of $A_{1}$ are called odd and an element that is either even or odd is said to be homogeneous.
Remark 3.4.
Equivalently, we can define a superring as a commutative monoid object with respect to the symmetric monoidal structure in the category $\mathrm{Ab}_{\mathbb{Z}_{2}}$.
Example 3.5.
Let $k$ be a field and $\{e_{1},...,e_{n}\}$ be a set. We can construct the free odd superalgebra by
(6) $\Lambda_{n}\coloneqq\bigwedge(e_{1},..,e_{n}).$
This algebra is called Graßmann algebra.
@@
A locally ringed superspace is a pair $X=(|X|,\mathcal{O}_{X})$, where $|X|$ is a topological space and $\mathcal{O}_{X}$ is a sheaf of superrings, such that for every point $x\in|X|$ $\mathcal{O}_{X,x}$ is a local superring.
Definition 3.7.
We define the category of affine superschemes to be
(7) $\mathrm{sAff}\coloneqq\mathrm{SRings}^{\mathrm{op}}.$
Definition 3.8.
A superscheme is a locally ringed superspace, which is locally isomorphic to the superspectrum of a superring.
@@
We can interpret the odd coordinates geometrically as fat points. Let $A$ be a superring, then the odd elements $\theta\in A_{1}$ satisfy the relation $\theta^{2}=0$. Let us now consider a superalgebra generated by a single degree 1 element over $\mathbb{C}$, i.e. $\mathbb{C[\theta]}$. We observe that if we forget the $\mathbb{Z}_{2}$-grading, there is an isomorphism
(8) $\mathbb{C[\theta]}\cong\mathbb{C}[\epsilon]/\epsilon^{2}.$
By this analogy, we can think of odd coordinates geometrically as thickening of the point or a first order Taylor expansion around the origin. Moreover, for a general superring $A$ we can also observe that, as topological spaces, we have a homeomorphism
(9) $\mathrm{Spec}(A)\cong\mathrm{Spec}(A/J).$
@@
Let $X$ be a superscheme, then we denote the sheaf generated by all degree 1 nilpotents by $J_{X}$. A superscheme X is called projected if the structural sequence
@@
In the category of supermanifolds, it turns out that every real supermanifold is graded in a non-natural way, due to Batchelor’s theorem [2]. So if we want to look for non-projected objects, we must look at complex supermanifolds.
@@
Let X be a superscheme of odd dimension m, then there exists a $J_{X}$-adic filtration of $\mathcal{O}_{X}$
@@
(11) $\mathrm{Gr}(\mathcal{O}_{X})\coloneqq J_{X}^{0}/J_{X}^{1}\oplus J_{X}^{1}/J_{X}^{2}\oplus...\oplus J_{X}^{m-1}/J_{X}^{m}\oplus J^{m}_{X}$
The graded superscheme associated to X is defined as
(12) $\mathrm{Gr}(X)\coloneqq(|X|,\mathrm{Gr}(\mathcal{O}_{X})).$
@@
A superscheme $X$ is called graded, if there exists an isomorphism $X\cong\mathrm{Gr}(X)$.
@@
The terminology in supergeometry is not yet set in stone. Some authors refer to this property as split [4], [32] and some authors call this property being a graded supermanifold [28]. We remind the reader to always carefully check, what is meant by the term graded or split when encountering supergeometry. It is also an unfortunate fact that all of these words are very overused in a multitude of contexts.
3.2. Algebraic Supergroups
Here we will introduce algebraic supergroups, following [20] and [21], the analogue of algebraic groups in supergeometry. While algebraic supergroups frequently arise in the context of supersymmetric physical models, their fundamental mathematical significance is perhaps best illustrated by Deligne’s theorem on tensor categories.
Theorem 3.15 ([8]).
Every $k$-tensor category $\mathcal{A}$ such that:
  1. (1)
    $k$ is an algebraically closed field of characteristic zero.
  2. (2)
    $\mathcal{A}$ is of subexponential growth.
is a neutral super Tannakian category, and there exists:
  1. (1)
    an affine algebraic supergroup $G$ whose algebra of functions $\mathcal{O}(G)$ is a finitely generated $k$-algebra.
  2. (2)
    a tensor-equivalence of categories
    $\mathcal{A}\simeq\operatorname{Rep}(G,\epsilon)$
    between $\mathcal{A}$ and the category of finite-dimensional representations of $G$.
This theorem roughly states that under mild conditions every tensor category can be realized as the category of representations of some algebraic supergroup.
In the standard literature, schemes are described as locally ringed spaces, whose structure sheaf encodes the algebraic data. Similarly, algebraic groups can be studied as locally ringed space. To make the group axioms compatible with the structure sheaf, we formulate the group axioms as a collection of commutative diagrams. This procedure results in the following